# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio Choice

### v5

@@ 2.1 Problem Formulation with State Variables and Neural Network Controls
We impose regularity conditions on the market parameters.
@@
Under Assumption 1, the SDE system for $(\mathbf{Y}_{t},X_{t})$ (where $X_{t}$ depends on the control processes $\bm{\pi}_{t}$ and $C_{t}$) admits a unique strong solution for sufficiently well-behaved controls. [^5] See Oksendal, (2013, Chs. 5–6) for the existence and uniqueness of solutions to SDEs.
The investor’s objective is to choose an admissible control pair $(\bm{\pi},C)$ that maximizes the expected discounted utility
[…]
subject to the wealth dynamics in Eq. (1) and the admissibility conditions stated later in Assumption 2. Here, $\delta>0$ is the subjective discount rate, and $\kappa\geq 0$ governs the strength of the bequest motive by weighting terminal utility.
Throughout, we assume preferences are represented by a utility function from the Hyperbolic Absolute Risk Aversion (HARA) family:
[…]
defined on the half-line $\mathcal{D}_{U}:=\{x>x_{\min}\}$, with $x_{\min}:=-a/b$. To ensure that the utility function is well-defined and the wealth process $X_{t}$ remains strictly positive, we assume $x_{\min}\geq 0$ throughout. This restriction precludes bankruptcy and reflects standard modeling practice in continuous-time portfolio choice under HARA preferences. This class includes the Constant Relative Risk Aversion (CRRA) utility for $a=0,b=1$, and the Constant Absolute Risk Aversion (CARA) utility for $\gamma=1,b=1$.
On its domain, $U$ is twice continuously differentiable, strictly increasing, and strictly concave. It also satisfies the Inada conditions. [^6] See Pennacchi, (2008, §3.2) for further discussion of the properties required of utility functions.:
$U^{\prime}(x)\to\infty\text{ as }x\downarrow x_{\min},\qquad U^{\prime}(x)\to 0\text{ as }x\to\infty.$
We consider control processes that are both economically sensible and mathematically tractable.
Assumption 2 (Admissible controls).
A control pair $(\bm{\pi},C)$ is admissible if the following conditions hold:
@@
[…]
@@
[…]
@@
    Domain-admissible wealth. Let $x_{\min}:=-a/b$ (as defined for the HARA utility). Then
[…]
@@
$\bm{\Omega}=\begin{pmatrix}\Psi&\rho\\[1.79997pt] \rho^{\top}&\Phi\end{pmatrix}$
@@
The assumptions are standard in continuous-time portfolio choice models. Specifically, Assumption 2(d) ensures that the wealth process remains within the domain of the utility function, i.e., $X_{t}>x_{\min}=-a/b$ almost surely, and also rules out arbitrage in this continuous-time setting (Dybvig and Huang, , 1988).
@@
Therefore, the objective functional (2) is well defined for every admissible control pair $(\bm{\pi},C)$.
The value function $V(t,x,\mathbf{y})$ associated with this problem represents the maximum expected utility achievable when starting from wealth $x$ and state $\mathbf{y}$ at time $t$. Formally,
[…]
where the supremum is taken over all admissible control processes on $[t,T]$, and $\mathbb{E}_{t,x,\mathbf{y}}[\cdot]:=\mathbb{E}[\cdot\mid X_{t}=x,\,\mathbf{Y}_{t}=\mathbf{y}]$ denotes the conditional expectation. The function $V$ is defined on $[0,T]\times(x_{\min},\infty)\times\mathbb{R}^{k}$.
@@
Assumption 4 ensures the diffusion matrix in the wealth dynamics is uniformly positive definite, which guarantees parabolic regularity of the associated Hamilton–Jacobi–Bellman (HJB) equation. We continue to work with the HARA utility function $U$ defined in Eq. (3), which is $C^{2}$, strictly increasing and concave on its domain $(x_{\min},\infty)$, and satisfies the Inada conditions. Together with the coefficient regularity in Assumption 1, these conditions yield the existence, uniqueness, and sufficient smoothness of the value function (e.g., $V\in C^{1,2,2}$); see Oksendal, (2013, Chs. 6–7).
In our numerical implementation, we parameterize the control policies using neural networks (NNs). Specifically, the portfolio allocation $\bm{\pi}_{t}$ and consumption rate $C_{t}$ are represented as functions of the current time $t$, wealth $X_{t}$, and state $\mathbf{Y}_{t}$, with functional forms determined by neural network parameters $\theta$ and $\phi$:
$\bm{\pi}_{t}=\bm{\pi}_{\theta}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},\qquad C_{t}=C_{\phi}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)}.$
The parameters $\theta$ and $\phi$ are optimized to maximize the objective functional $J$ defined in Eq. (2).
2.2 Optimality Conditions via Pontryagin’s Maximum Principle
We now apply Pontryagin’s Maximum Principle (PMP) to the continuous-time portfolio choice problem formulated in Section 2.1, where market parameters depend on the exogenous state process $\{\mathbf{Y}_{t}\}_{0\leq t\leq T}$. PMP provides necessary conditions for optimality via a system of forward-backward stochastic differential equations (FBSDEs), which couple the dynamics of the state variables with those of the associated costate (or adjoint) processes.
The system state consists of the investor’s wealth $X_{t}$ and the exogenous state variables $\mathbf{Y}_{t}\in\mathbb{R}^{k}$. The corresponding costate processes represent the marginal value of state variables with respect to the value function $V$: the scalar process $\lambda_{t}:=V_{x}(t,X_{t},\mathbf{Y}_{t})$ corresponds to the sensitivity to wealth, and the vector process $\bm{\eta}_{t}:=V_{\mathbf{y}}(t,X_{t},\mathbf{Y}_{t})\in\mathbb{R}^{k}$ captures sensitivity to the exogenous state.
Under an optimal control policy $(\bm{\pi}_{t}^{*},C_{t}^{*})$, the forward state dynamics are governed by:
with initial conditions $X_{0}^{*}=x_{0}>x_{\min}$ and $\mathbf{Y}_{0}=\mathbf{y}_{0}\in\mathbb{R}^{k}$. Recall that $\bm{\sigma}(t,\mathbf{Y}_{t})$ is the asset volatility matrix, and $\bm{\Omega}$ is the covariance matrix of the combined Brownian motion $\mathbf{W}_{t}=\bigl{(}\mathbf{W}^{X\top}_{t},\,\mathbf{W}^{Y\top}_{t}\bigr{)}^{\top}$.
To characterize optimal controls, PMP introduces a Hamiltonian function that incorporates both the instantaneous utility and the effects of the controls on the system’s drift and diffusion. In problems with control-dependent diffusion, the Hamiltonian must also account for the interaction with costate diffusion terms (i.e., the $\mathbf{Z}$-components of the adjoint BSDEs). Following standard stochastic control formulations (e.g., Yong and Zhou, , 2012), the Hamiltonian $\mathcal{H}$ is defined as:
$\mathcal{H}=e^{-\delta t}\,U(C)+\lambda\left[x\bigl{(}r+\bm{\pi}^{\top}(\bm{\mu}-r\mathbf{1})\bigr{)}-C\right]+\bm{\eta}^{\top}\bm{\mu}_{Y}+\mathbf{Z}^{\lambda X}\cdot(x\,\bm{\sigma}^{\top}\bm{\pi})+(\mathbf{Z}^{\eta Y})^{\top}\bm{\sigma}_{Y},$ (6)
where $\lambda$ and $\bm{\eta}$ are the costate processes, $\mathbf{Z}^{\lambda X}$, $\mathbf{Z}^{\eta X}$, and $\mathbf{Z}^{\eta Y}$ represent the diffusion coefficients of the backward SDEs associated with $\lambda_{t}$ and $\bm{\eta}_{t}$, and all functions are evaluated at $(t,x,\mathbf{y},\bm{\pi},C)$, though we suppress this for clarity.
Complementing the forward dynamics, the optimal costate processes $(\lambda_{t}^{*},\bm{\eta}_{t}^{*})$ satisfy a system of backward stochastic differential equations (BSDEs), driven by the partial derivatives of the Hamiltonian $\mathcal{H}$ with respect to the state variables:
$-d\bm{\eta}_{t}^{*}$ $=\mathcal{H}_{\mathbf{y}}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*}\bigr{)}\,dt-\mathbf{Z}_{t}^{\eta X*}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\eta Y*}\,d\mathbf{W}_{t}^{Y}.$ (8)
Here, $\mathcal{H}_{x}:=\partial\mathcal{H}/\partial x$ and $\mathcal{H}_{\mathbf{y}}:=\nabla_{\mathbf{y}}\mathcal{H}$ denote the partial derivatives of the Hamiltonian with respect to wealth and the state vector, respectively.
The terms $\mathbf{Z}_{t}^{\lambda X}\in\mathbb{R}^{1\times n}$, $\mathbf{Z}_{t}^{\lambda Y}\in\mathbb{R}^{1\times k}$, $\mathbf{Z}_{t}^{\eta X}\in\mathbb{R}^{k\times n}$, and $\mathbf{Z}_{t}^{\eta Y}\in\mathbb{R}^{k\times k}$ are adapted processes. [^7] These terms arise from the martingale representation theorem applied to the backward dynamics of the costate variables. See Karatzas and Shreve, (1991) for a detailed treatment of the martingale representation theorem. These processes are collectively denoted by $\mathbf{Z}_{t}$.
Note that $\mathcal{H}_{x}$ and $\mathcal{H}_{\mathbf{y}}$ generally depend not only on the state, costate, and control variables, but also on the $\mathbf{Z}_{t}$ terms themselves. This reflects the coupling between the forward and backward systems, which arises from the second-order sensitivity of the value function.
The terminal conditions for the BSDEs are determined by the structure of the objective functional:
@@ 2.2 Optimality Conditions via Pontryagin’s Maximum Principle
In addition to Assumptions 1, 2, and 4, we assume:
@@
    $C^{1}$-regularity with respect to $\mathbf{y}$. All coefficient functions—$\bm{\mu}_{Y},\,\bm{\sigma}_{Y},\,\bm{\mu},\,\bm{\sigma},\,r$—are continuously differentiable with respect to the state vector $\mathbf{y}$. Moreover, their partial derivatives $\partial_{\mathbf{y}}f$ are globally Lipschitz continuous in $\mathbf{y}$ and satisfy linear growth bounds. These conditions ensure that the drift terms in the costate BSDEs, including expressions such as $\mathcal{H}_{\mathbf{y}}$, satisfy the regularity requirements for the existence and uniqueness of solutions; see Yong and Zhou, (2012, Chapter 2 and 6).
The core requirement of Pontryagin’s Maximum Principle is that the optimal controls $(\bm{\pi}_{t}^{*},C_{t}^{*})$ maximize the Hamiltonian $\mathcal{H}$ (6) pointwise in time, almost surely, given the optimal state and costate processes $(X_{t}^{*},\mathbf{Y}_{t},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*})$. Assuming interior solutions allows us to derive the first-order conditions (FOCs) for optimality. Maximizing with respect to the consumption rate $C_{t}$ yields the standard FOC:
$e^{-\delta t}\,U^{\prime}(C_{t}^{*})=\lambda_{t}^{*}.$ (9)
This classical result links optimal consumption to the marginal value of wealth via the inverse of the marginal utility: $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$.
For the portfolio weights $\bm{\pi}_{t}$, we differentiate $\mathcal{H}$ with respect to $\bm{\pi}$, accounting for both the drift and diffusion terms that depend on the control. Setting the gradient $\nabla_{\bm{\pi}}\mathcal{H}=\mathbf{0}$, we obtain:

### v6

@@ Appendix A Regularity Conditions and Assumptions for a General Multi-Asset Model
We impose regularity conditions on the market parameters of the general multi-asset model in Section 3.1.
@@
Under Assumption 1, the SDE system for $(\mathbf{Y}_{t},X_{t})$ (where $X_{t}$ depends on the control processes $\bm{\pi}_{t}$ and $C_{t}$) admits a unique strong solution for sufficiently well-behaved controls. For the existence and uniqueness of the solutions to SDEs, see chapters 5 and 6 in Oksendal, (2013).
Now we define the admissible set $\mathcal{A}(x,\mathbf{y})$ for our stochastic control problem.
Definition 1 (Admissible set).
A pair $(\bm{\pi},C)$ is admissible if the following conditions hold:
@@
    $\mathbb{E}\!\int_{0}^{T}\left(\lVert\bm{\pi}_{t}\rVert^{2}+C_{t}\right)\,dt\;<\;\infty.$
@@
    $\lVert\bm{\pi}_{t}\rVert\;<\;\Pi_{\max},\qquad 0\;\leq\;C_{t}\;<\;\xi\,X_{t},\qquad\forall t\in[0,T]\text{ a.s.}$
@@
    Doadmissible wealth. Let $x_{\min}:=-a/b$ (as defined for the HARA utility). Then
    $X_{t}\;>\;x_{\min},\qquad\forall t\in[0,T]\text{ a.s.}$
@@
$\bm{\Omega}=\begin{pmatrix}\Psi&\rho\\[1.84995pt] \rho^{\top}&\Phi\end{pmatrix}$
@@
The definition of an admissible set is standard in continuous-time portfolio choice models. Specifically, the doadmissible wealth condition in (d) ensures that the wealth process remains within the domain of the utility function, i.e., $X_{t}>x_{\min}=-a/b$ almost surely, and also rules out arbitrage in this continuous-time setting (Dybvig and Huang, , 1988).
@@
Therefore, the value function $V(0,x,\mathbf{y})$ is well-defined for every admissible control pair $(\bm{\pi},C)$. We set all admissible pairs by the admissible set $\mathcal{A}(x,y)$.
@@
Assumption 3 ensures the diffusion matrix in the wealth dynamics is uniformly positive definite, which guarantees parabolic regularity of the associated Hamilton–Jacobi–Bellman (HJB) equation. We continue to work with the HARA utility function $U$ defined in Eq. (7), which is $C^{2}$, strictly increasing and concave on its domain $(x_{\min},\infty)$, and satisfies the Inada conditions. Together with the coefficient regularity in Assumption 1, these conditions yield the existence, uniqueness, and sufficient smoothness of the value function (e.g., $V\in C^{1,2,2}$); see Oksendal, (2013, Chapter 6 and 7).
@@
In addition to Assumptions 1, 3, and an addmissible set in Definition 1, we assume: